Two 10-cm-diameter charged rings face each other, 20 cm apart. The left ring is charged to -15 nC and the right ring is charged to +15 nC

You may want to review (Pages 641-643).


Part A

What is the magnitude of the electric field \vec{E} at the midpoint between the two rings?

Part B

What is the direction of the electric field \vec{E} at the midpoint between the two rings?

Part C

What is the magnitude of the force on a proton at the midpoint?

Part D

What is the direction of the force \vec{F} on a proton at the midpoint?
Answer #1

a)

electric field at mid point

E = (q/A) / eo

E = 15* 10^-9 / ( 3.14* 0.05^2* 8.85* 10^-12)

E = 2.16* 10^-5 N/C

=====

b)

field directed to the left

=====

c)

now force at mid point of the system

F = q E = 1.6* 10^-19* 15* 10^-9 / (3.14* 0.05^2* 8.85* 10^-12)

F = 3.45* 10^-14 N

=====

d)

force directed towards the left plate

=====

comment before rate in case any doubt, will reply for sure.. goodluck